A partnership has the following account balances: Cash, $82,000; Other Assets, $600,000; Liabilities, $235,000; Nixon (50 percent of profit and losses), $210,000; Cleveland (30 percent), $145,000; Pierce (20 percent), $92,000. The company liquidates, and $19,500 becomes available to the partners. Who gets the $19,500? Determine how much should be distributed to each partner?
Answer #1

Total Capital = 210000 + 145000 + 92000 = 447000

Maximum Potential Loss = Total Capital Available - Capital Available to Partners = 447000 - 19500 = $427,500

Nixon Cleveland Pierce
Reported Balances (1) $210000 $145000 $92000
Split of Anticipated Loss of $427,500 in the ratio of 5:3:2 (2) 213,750 128,250 85,500
Potential Balances (1-2) (3,750) 16,750 6,500
Nixon's Deficit (Split in the ratio of 3:2) (3) 3,750 2,250 1,500
Current Cash Distribution (Potential Balances - 3) 0 14,500 5,000